OpenStudy (grandmoffchris):

-2/9 x (-3/14) = Will fan and medal, Please help!!

OpenStudy (anonymous):

okay are you with me

OpenStudy (grandmoffchris):

@imqwerty @ChantySquirrel1129** @AloneS

OpenStudy (grandmoffchris):

i am

OpenStudy (anonymous):

-2 * -3

OpenStudy (grandmoffchris):

lets do this

OpenStudy (anonymous):

alright

OpenStudy (grandmoffchris):

-2 x -3?

OpenStudy (anonymous):

yes

OpenStudy (grandmoffchris):

so i multiply and get answer?

OpenStudy (anonymous):

-2 * -3

OpenStudy (grandmoffchris):

although didnt put what it said, Multiply. Express your answer in lowest terms.

OpenStudy (gabbyalicorn):

Is -2/9 & -3/14 shown as a fraction or are they being divided?

OpenStudy (grandmoffchris):

Ah, umm

OpenStudy (grandmoffchris):

fraction

OpenStudy (anonymous):

yes

OpenStudy (grandmoffchris):

okay i gtg,

OpenStudy (anonymous):

start up with dividing 3 in 9 and -3

OpenStudy (grandmoffchris):

im in k12, and my class starts at 9:30 am

OpenStudy (grandmoffchris):

although ill brb actually

OpenStudy (grandmoffchris):

at 10

OpenStudy (gabbyalicorn):

−2/9(−3/14) =(−2/9)(−3/14) =(−2)(−3)/(9)(14) =6/126 =1/21(Decimal: 0.047619) This is what I got mine starts at 1 xD

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

common you are meant to teach him not solve it

OpenStudy (grandmoffchris):

what

OpenStudy (gabbyalicorn):

I showed him my steps. Some people are visual learners. xD

OpenStudy (grandmoffchris):

daniel is right tho

OpenStudy (grandmoffchris):

Oh lol

OpenStudy (grandmoffchris):

that wasnt answer @gabbyalicorn

OpenStudy (grandmoffchris):

@eggman

OpenStudy (mathmale):

-2/9 x (-3/14) = Will fan and medal, Please help!! I am assuming that you mean the following:\[(-2/9) (-3/14) = ? \]